Chapter 9 Problem 1.7 Fill in the missing values in the following table of data collected...

  1. Chapter 9 Problem 1.7

Fill in the missing values in the following table of data collected in the household survey for February 2017:

Working-Age Population

Employment

152,528,000

Unemployment

?

Unemployment Rate

4.7%

Labor force

?

Labor force participation rate

63.0%

Employment-population ratio

?

Source: U.S. Department of Labor, Bureau of Labor Statistics, The Employment Situation – February 2017, March 10, 2017.
Answer #1

Unemployment rate = unemployed/(unemployed+employed)

0.047 = Unemployed/(Unemployed+Employed)

0.047 = X/(152528000+X)

0.047*(152528000+X) =X

0.047X+7168816 =X

7168816 = X-0.047X

X = 7168816/0.953 = 7522367.26

a) Unemployed = 7522367

b) Labor force = unemployed+employed 7522367+152528000 = 160050367

Now, since labor force participation rate = 0.63

LFPR = Labor force/Working age population

Let working age population be X

0.63 = 160050367/X

X = 254048202

C) Employment population ratio = Employment/Working age population*100

= 15252800/254048202*100

= 60.04%